OpenStudy (anonymous):

The temperature was 56 degrees this morning but dropped one degree per hour through the rest of the day. Read over problem and look at the graph you made to decide how to answer this question. Would it make sense to predict the temperature in 45 hours? Why or why not (required).

OpenStudy (anonymous):

hint: 1. John began his job making $25 the first day. After that he was paid $6.75 per hour. Write the linear equation for this problem in y=mx+b form ans...-(1.) y = 6.75x + 25

OpenStudy (campbell_st):

the problem with the question and a linear, straight line graph is that it doesn't allow or temperature rises.... temp intercept = 56, gradient = -1 (dropping 1 degree per hour) E.g the linear equation for you problem is t = -h + 56 in your question the prediction would be t = -45 + 56 t = 11 the graph is quadrants 1 and 4 as time can't be negative.